OpenStudy (anonymous):

Find g(1) if g(x) = x 2 + 1.

OpenStudy (anonymous):

anybody ?

OpenStudy (nurali):

g(x) = x^2 + 1. x=1 g(1)=(1)^2+1 g(1)=1+1 g(1)=2

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

Anytime.